# CoRanking: Collaborative Ranking with Small and Large Ranking Agents

### v2

@@ 3.2 Overall Results
The main experimental results are presented in Table 1, from which we draw the following observations:
(1) Our collaborative reranking framework CoRanking achieves the best average performance (51.39) across all datasets, outperforming both single rerankers and other collaborative baselines. Notably, it even surpasses the LLR by about 0.5 points on average. These results demonstrate that our framework could effectively align the passage order generated by the SLR (3B) with the preferences of the large language model (LLR), thereby achieving better ranking performance than only using LLM reranker LLR.
(2) Regarding single rerankers, our SLR (3B) achieves an average score of 47.90, outperforming all the pointwise rerankers including monoBERT, monoT5, RankT5 and Qwen-pointwise, and even surpassing 7B-scale RankVicuna and RankZephyr. Furthermore, SLR surpasses Qwen-pointwise by about 1.8 average points under identical training conditions, proving listwise reranking’s superiority through comparison of multiple passages for precise relevance assessment.
(3) Collaborative reranker baselines (e.g., RankT5 (3B) + LLR) improve over corresponding single rerankers (e.g.RankT5 (3B)). However, these baselines still lags behind and LLR and our CoRanking, indicating that directly feeding small rerankers’ ranking results to LLR fails to align with LLR’s order preference.
@@ 3.3 Ablation Study
First, to evaluate our proposed $\text{S}^{3}$ sampling strategy for generating the preference pairs of DPO, we introduce a variant, denoted as “w/o $\text{S}^{3}$”. This variant uses random sampling for preference pair selection, without ensuring the significant reward of winner prediction and the reward gap between the winner and loser prediction. The performance degradation across all benchmarks (e.g., 1.54-point drop on TREC Avg.) demonstrate the effectiveness of our $\text{S}^{3}$ sampling in identifying valuable preference pairs for effective DPO training.
We further remove our POA to validate its alignment effectiveness, denoted as “w/o POA”. In this case, the LLR directly reranks the top-20 passages after SLR’s reranking. The notable performance decline (especially 1.6-point decrease on BRIGHT Avg.) demonstrates that POA effectively align with the passage order preference of LLR.
Finally, to validate our human-label-enhanced ranking construction (HRC) for fine-tuning SLR, we create a variant (w/o HRC) which generates ranking label by utilizing LLR to rerank training passages only one time and choosing not to use human label for filtering high-quality ranked lists. The significant performance drops across benchmarks (especially BEIR) highlight HRC’s effectiveness on generating high-quality ranking labels.
3.4 Efficiency Analysis
@@ 3.5 Detailed Analysis of POA
Our POA is designed to adjust the order of top-20 passages reranked by SLR to align them more with the preferences of the LLR. However, whether this adjustment improves or degrades the ranking quality of top-20 passages remains unknown. We believe this is an interesting question because it helps us understand the reasons why POA works.
In this part, we selected four datasets—DL19, Robust04, News, and Economics—and reported the changes in the ranking metric NDCG@10 after ranking by SLR, POA, and LLR respectively. We also report the performance of the ablated variant SLR+LLR (i.e., directly using LLR to rerank top-20 passages from SLR without POA’s adjustment) as a baseline. The results are shown in Figure 4. After POA’s adjustment, the ranking metric on three datasets—DL19, Robust04, and News—decreases (except for Economics dataset). This indicates that a passage ranking list with high-ranking metric does not necessarily align with LLR’s preferences. This may be due to the positional bias of input tokens during the pre-training of LLMs, as revealed in previous studies Tang et al. (2024); Liu et al. (2024a).
@@ 3.6 Generalization on Other LLMs
In the previous experiments, we use Qwen2.5-72B-Instruct as our LLR model. However, it remains unclear whether our proposed CoRanking framework can generalize effectively to other LLR models. To evaluate this, we conducted additional experiments with two different LLRs: GPT-4o [^3] The version of GPT-4o we used is gpt-4o-2024-08-06. and DeepSeek-V3, and compare our CoRanking with three collaborative rerankers (“RankT5 (3B) + LLR”, “Qwen-pointwise (3B) + LLR”, and “SLR + LLR”). Note that “SLR + LLR” represents using LLR to directly rerank the top-20 passages reranked by SLR without POA. For evaluation, we select one dataset from each of the three benchmarks: DL19 from TREC, News from BEIR, and Earth Science from BRIGHT. As illustrated in Table 3, our CoRanking consistently outperforms the best baseline “SLR + LLR” by about 1 point on average, regardless of the LLR used. This demonstrates the positional preference alignment of our POA could generalize to different LLMs.
@@ 4.1 LLMs for Passage Ranking
The application of large language models (LLMs) into retrieval systems Zhu et al. (2023) has driven substantial research efforts in passage ranking tasks. Current LLM-based ranking methods can be categorized into three categories: pointwise, pairwise, and listwise. Pointwise methods Liang et al. (2022); Sachan et al. (2022); Zhuang et al. (2023a); Liu et al. (2024e, d) compute the relevance score for each query-passage pair. Pairwise methods Qin et al. (2023); Luo et al. (2024) compare two passages at a time and identify the more relevant one. Listwise methods Sun et al. (2023); Pradeep et al. (2023a); Reddy et al. (2024); Liu et al. (2024c, b); Chen et al. (2024) directly take a list of passages as the input and output the passage identifiers in descending order of their relevance. Through comparing multiple passages, listwise reranker could assess the relevance more comprehensively and yield better ranking results. Due to the superior performance of listwise ranking Sun et al. (2023), it has attracted the interest of many researchers. For example,  Tang et al. (2024) propose a permutation self-consistency method to address the sensitivity of listwise ranking to initial passage order and achieve a more accurate and positionally unbiased ranking. Besides,  Pradeep et al. (2023a) and  Pradeep et al. (2023b) propose to distill the knowledge of strong listwise reranker (e.g., GPT-4) into moderate-sized models to enhance ranking efficiency while reducing computational overhead.
The superior performance of LLM listwise rerankers is often attributed to the utilization of large-scale LLMs and the sliding window strategy, which leads to efficiency issue in ranking. In this paper, we introduce a novel framework that leverages the collaborative ranking of large-size rerankers and small-size rerankers to enhance the ranking efficiency while maintaining ranking effectiveness.
@@ 4.2 Preference Alignment
Preference alignment has emerged as a critical research direction, aiming to align model behaviors with human or task-specific preferences. Early approaches primarily relied on reinforcement learning from human feedback (RLHF), where models are fine-tuned using reward signals derived from human preference annotations. To reduce the computational complexity of RLHF, recent studies proposed more efficient alternatives. Direct Preference Optimization (DPO) Rafailov et al. (2023) reformulates preference learning as a supervised objective, bypassing explicit reward modeling while achieving competitive performance. Recent extensions like Step-DPO Lai et al. (2024) further decompose preference alignment into multi-step optimization. Besides, Dong et al. (2023) introduced rejection sampling fine-tuning to align LLMs with human preferences without requiring reinforcement learning pipelines. In this paper, to address the positional sensitivity of LLM listwise reranker, we propose training a passage order adjuster with the DPO algorithm to align the passage order generated by the small reranker with the positional preferences of the LLM-based listwise reranker.
@@ 5 Conclusion
In this paper, we introduce a novel collaborative ranking framework CoRanking which combines small and large rerankers for more efficient and effective passage ranking. Specifically, we first train a small listwise reranker SLR to move the relevant passages to the top of the list. Then we introduce a passage order adjuster POA to reorder the top-ranked passages to better align with the positional preference of LLM listwise reranker. Finally, the LLM listwise reranker is utilized to solely rerank the top passages. Extensive experiments demonstrate the efficiency advantage and superior performance of CoRanking.
@@
We acknowledge some limitations in our work. Firstly, our framework consists of three models, making its online deployment more complex compared to using a single unified reranker. Secondly, we only use Qwen2.5-3B-Instruct as the backbone model of our SLR and POA. We believe it is meaningful to attempt using different backbones to demonstrate the effectiveness of our CoRanking framework.
@@ Appendix B Datasets Details
Appendix C Baselines Details
$\bullet$ LLR (72B): Our teacher model that employs a zero-shot sliding window strategy for passage re-ranking.

### v3

@@ 3.2 Overall Results
The main experimental results are presented in Table 1, from which we could draw the following observations:
(1) Our collaborative reranking framework CoRanking achieves the best average performance (51.39) across all datasets, outperforming both single rerankers and other collaborative baselines. Notably, it even surpasses the large listwise reranker LLR by about 0.5 points on average. These results demonstrate that our framework could effectively align the passage order generated by the SLR (3B) with the preferences of the large language model (LLR), thereby achieving better ranking performance than only using LLR. (2) Regarding single rerankers, our SLR (3B) achieves an average score of 47.90, outperforming all the pointwise rerankers, including monoBERT, monoT5, RankT5, and Qwen-pointwise, and even surpassing 7B-scale RankVicuna and RankZephyr. Furthermore, SLR surpasses Qwen-pointwise by about 1.8 average points under identical training conditions, proving listwise reranking’s superiority through comparison of multiple passages for precise relevance assessment.
(3) Collaborative reranker baselines (e.g., RankT5 (3B) + LLR) improve over corresponding single rerankers (e.g., RankT5 (3B)). However, these baselines still lag behind LLR and our CoRanking, indicating that directly feeding small rerankers’ ranking results to LLR fails to align with LLR’s order preference.
3.3 Efficiency Analysis
@@ 3.4 Ablation Study
First, to evaluate our proposed $\text{S}^{3}$ sampling strategy for generating the preference pairs of DPO, we introduce a variant, denoted as “w/o $\text{S}^{3}$”. This variant uses random sampling for preference pair selection, without ensuring the significant reward of winner prediction and the reward gap between the winner and loser predictions. The performance degradation across all benchmarks (e.g., 1.54-point drop on TREC Avg.) demonstrates the effectiveness of our $\text{S}^{3}$ sampling in identifying valuable preference pairs for effective DPO training.
We further remove our POA to validate its alignment effectiveness, denoted as “w/o POA”. In this case, the LLR directly reranks the top-20 passages after SLR’s reranking (i.e., the naive collaboration mentioned in Figure 1 (B)). The notable performance decline (especially 1.6-point decrease on BRIGHT Avg.) demonstrates that POA could effectively align with the order preference of LLR.
Finally, to validate our human-label-enhanced ranking construction (HRC) for fine-tuning SLR, we create a variant (w/o HRC) which generates ranking labels by utilizing LLR to rerank training passages only once and choosing not to use human labels for filtering high-quality ranked lists. The significant performance drops across benchmarks (especially BEIR) highlight HRC’s effectiveness in generating high-quality ranking labels.
@@ 3.5 Detailed Analysis of POA
Our POA is designed to adjust the order of the top-20 passages reranked by SLR to align them more with the preferences of the LLR. However, whether this adjustment improves or degrades the ranking quality of top-20 passages remains unknown. We believe this is an interesting question because it helps us understand the reasons why POA works.
In this part, we selected four datasets—DL19, Robust04, News, and Economics—and reported the changes in the ranking metric NDCG@10 after ranking by SLR, POA, and LLR, respectively. We also report the performance of the naive collaboration SLR+LLR as a baseline. The results are shown in Figure 4. After POA’s adjustment, the ranking metric on three datasets—DL19, Robust04, and News—decreases (except for the Economics dataset). This indicates that a passage ranking list with a high-ranking metric does not necessarily align with LLR’s preferences. This may be due to the positional bias of input tokens during the pre-training of LLMs Tang et al. (2024). Moreover, due to the stronger ranking capability, LLR further improves effectiveness compared to SLR and SLR+POA, demonstrating its necessity in the CoRanking framework.
@@ 3.6 Generalization on Other LLMs
In the previous experiments, we use Qwen2.5-72B-Instruct as our LLR model. However, it remains unclear whether our proposed CoRanking framework can generalize effectively to other LLR models. To evaluate this, we conduct additional experiments with two different LLRs: GPT-4o [^4] The version of GPT-4o we used is gpt-4o-2024-08-06. and DeepSeek-V3, and compare our CoRanking with three collaborative rerankers (“RankT5 (3B) + LLR”, “Qwen-pointwise (3B) + LLR”, and “SLR + LLR”). Note that “SLR + LLR” represents the naive collaboration in Figure 1 (B). For evaluation, we select one dataset from each of the three benchmarks: DL19 from TREC, News from BEIR, and Earth Science from BRIGHT. As illustrated in Table 3, our CoRanking consistently outperforms the best baseline “SLR + LLR” by about 1 point on average, regardless of the LLR used. This demonstrates that the order preference alignment of our POA could generalize to different LLMs.
@@ 4.1 LLMs for Passage Ranking
The application of large language models (LLMs) into retrieval systems Zhu et al. (2023) has driven substantial research efforts in passage ranking tasks. Current LLM-based ranking methods can be categorized into three categories: pointwise, pairwise, and listwise. Pointwise methods Liang et al. (2022); Sachan et al. (2022); Zhuang et al. (2023a); Liu et al. (2024e, d) compute the relevance score for each query-passage pair. Pairwise methods Qin et al. (2023); Luo et al. (2024) compare two passages at a time and identify the more relevant one. Listwise methods Sun et al. (2023); Pradeep et al. (2023a); Reddy et al. (2024); Liu et al. (2024c, b); Chen et al. (2024) directly take a list of passages as the input and output the passage identifiers in descending order of their relevance. Through comparing multiple passages, a listwise reranker could assess the relevance more comprehensively and yield better ranking results. Due to the superior performance of listwise ranking, it has attracted the interest of many researchers Tang et al. (2024); Pradeep et al. (2023b); Liu et al. (2025); Fan et al. (2025).
The superior performance of LLM listwise rerankers is often attributed to the utilization of large-scale LLMs and the sliding window strategy, which leads to an efficiency issue in ranking. In this paper, we introduce a novel framework that leverages the collaborative ranking of large rerankers and small rerankers to enhance the ranking efficiency while maintaining ranking effectiveness.
@@ 4.2 Preference Alignment
Preference alignment has emerged as a critical research direction, aiming to align model behaviors with human or task-specific preferences. Early approaches primarily relied on reinforcement learning from human feedback (RLHF), where models are fine-tuned using reward signals derived from human preference annotations. To reduce the computational complexity of RLHF, recent studies propose more efficient alternatives, such as DPO Rafailov et al. (2023) and Step-DPO Lai et al. (2024). Direct preference optimization (DPO) Rafailov et al. (2023) reformulates preference learning as a supervised objective, bypassing explicit reward modeling while achieving competitive performance. In this paper, to address the positional sensitivity of a large listwise reranker, we propose training a passage order adjuster with the DPO algorithm to align the passage order generated by the small listwise reranker with the order preference of the large listwise reranker.
@@ 5 Conclusion
In this paper, we introduce a novel collaborative ranking framework, namely CoRanking, to combine small and large listwise rerankers for more efficient and effective passage ranking. Specifically, we first train a small listwise reranker SLR to move the relevant passages to the top of the list. Then we introduce a passage order adjuster POA to rerank the top-ranked passages to better align with the order preference of a large listwise reranker. Finally, the large listwise reranker is utilized to solely rerank the top passages. Extensive experiments demonstrate the efficiency advantage and superior performance of our CoRanking.
@@
We acknowledge some limitations in our work. First, our framework consists of three models, making its online deployment more complex compared to using a single unified reranker. Second, we only use Qwen2.5-3B-Instruct as the backbone model of our SLR and POA. We believe it is meaningful to attempt using different backbones (such as Llama-series models) and models of varying sizes (such as Qwen2.5-1.5B-Instruct) to further demonstrate the generalization of our CoRanking framework.
Acknowledgments
This work was supported by National Natural Science Foundation of China No. 62272467, Beijing Municipal Science and Technology Project No. Z231100010323009, Beijing Natural Science Foundation No. L233008. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
@@ Appendix B Datasets Details
Appendix C Baselines Details
$\bullet$ $\textbf{RankMistral}_{\text{100}}$ Liu et al. (2024c): $\textbf{RankMistral}_{\text{100}}$ is a full reranker trained using a full ranking list generated by GPT-4o. It directly takes all the passages as input and output the reranked list without relying on a sliding window strategy.
$\bullet$ LLR (72B): The large listwise reranker that employs a zero-shot sliding window strategy for passage re-ranking.
Appendix D More Implementation Details
Unlike previous studies Pradeep et al. (2023b); Liu et al. (2024c) that adopted GPT-series models (e.g., GPT-4o) as the teacher model, we select Qwen2.5-72B-Instruct as it demonstrates comparable and even better performance to GPT-4o across most datasets while being more cost-effective. Nevertheless, we have included additional experimental results using GPT-4o as the LLR model in Section 3.6. We train the SLR for 4 epochs with a learning rate of 5e-6 and a batch size of 1. The batch size, learning rate, and the hyperparameter $\beta$ of DPO are set as 1, 1e-6, and 0.4, respectively.